OpenStudy (anonymous):

can i get some quick help with 3 geometry questions??

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

what are the questions?

OpenStudy (anonymous):

okay here is the first one,

OpenStudy (anonymous):

i dont have a picture of the triangle, how do you want me to explain that?

OpenStudy (anonymous):

A source of laser light sends rays AB and AC toward two opposite walls of a hall. The light rays strike the walls at points B and C, as shown below: A source of laser light is at point A on the ground between two parallel walls. The walls are perpendicular to the ground. AB is a ray of light which strikes the wall on the left at point B. The length of AB is 40m. AC is a ray of light which strikes the wall on the right at point C which is 80 m above the ground. The ray AB makes an angle of 60 degrees with the ground. The ray AC makes an angle of 45 degrees with the ground. What is the distance between the walls? 80 m 180 m 90 m 100 m

OpenStudy (anonymous):

use the draw function and draw it

OpenStudy (anonymous):

thats not the right question lol

OpenStudy (anonymous):

dont pay attention to that haha

OpenStudy (anonymous):

ok haha

OpenStudy (anonymous):

what is the value of sin x ? 

OpenStudy (anonymous):

like i know pretty much what i am doing, just i dont have a brain for math lol. so im a a little slower getting started on questions :/

OpenStudy (anonymous):

you want to pick sin, cos, or tan and use the sides to find x, so for examples you could use sin(x)=12/13 which then equals 67.38

OpenStudy (anonymous):

okay so

OpenStudy (anonymous):

im trying to comprehend this lol

OpenStudy (anonymous):

do you understand how to use sin, cos, and tan?

OpenStudy (anonymous):

so you would use sin?

OpenStudy (anonymous):

yes i do, somewhat lol

OpenStudy (anonymous):

sin = opposite over hypotenuse

OpenStudy (anonymous):

Yes that is correct! that numbers would you plug in for sin?

OpenStudy (anonymous):

*what numbers, not that number

OpenStudy (anonymous):

13 is my hopotenuse. 13 is my opposite?

OpenStudy (anonymous):

i mean 12

OpenStudy (anonymous):

Yup! Good job

OpenStudy (anonymous):

then heres my problem , i forget what to do next :(

OpenStudy (anonymous):

ok, so....\[\sin x=\frac{ opposite }{ hypotenuse }\]\[\sin x=\frac{ 12 }{ 13}\]\[x=\sin^{-1} \frac{ 12 }{ 13 }\] x=67.38

OpenStudy (anonymous):

so you divided?

OpenStudy (anonymous):

anything else?

OpenStudy (anonymous):

yeah i have two more questions, but just to be clear. i think im confusing myself.

OpenStudy (anonymous):

what did you do with 12/13 to get that answer

OpenStudy (anonymous):

you need to take the inverse of sin on both sides to get x alone, so you would do Sin^-1(12/13)

OpenStudy (anonymous):

im going to open up the next question so i can medal you again

OpenStudy (anonymous):

@GWiz

OpenStudy (anonymous):

ok